OpenStudy (anonymous):

Please help! How much pure antifreeze must be added to 12 gallons of 20% antifreeze to make a 80% antifreeze solution ?

OpenStudy (anonymous):

0.2 x 12 = 2.4 we have 2.4 gallons of pure antifreeze in the solution at the moment if we add x gallons we will have (x+2.4) pure antifreeze in solution which has total volume of (x+12) so we want to find x in this equation:\[\frac{x+2.4}{x+12} = \frac{80}{100}\]

OpenStudy (anonymous):

do you see how i got that?

OpenStudy (anonymous):

\[100(x+2.4) = 80(x+12)\]

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

\[100x + 240 = 80x + 960\]\[20x = 720\] so yes:\[x = 36\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem